OpenStudy (anonymous):

How do you solve 9x^2-30x+25=0

OpenStudy (campbell_st):

well its a perfect square quadratic... can you factorise it..?

OpenStudy (anonymous):

This is what I have so far 9x^2=4.5x 25 is 52ab = 2(x)(5)=10x I dont think im doing this right ?

OpenStudy (campbell_st):

look at it this way \[(3x)^2 = 9x^2\]

OpenStudy (anonymous):

so how do I solve it with (3x^2)

OpenStudy (campbell_st):

please read the information \[3x \times 3x = 9x^2.... or .....(3x)^2 = 9x^2\] 25 = 5 x 5 middle term is negative so its -5 now look at the middle term -30x is -30x/2 = 3x * -5...?

OpenStudy (anonymous):

I still do not know how to solve this.

OpenStudy (campbell_st):

ok... are you happy about 3x * 3x = 9x^2 and 25 = -5 * -5

OpenStudy (anonymous):

yes i get that, but now what do you do or is that it?

OpenStudy (campbell_st):

ok so the factorised form is \[9x^2 - 30x + 25 = (3x - 5)^2\] so \[(3x - 5)^2 = 0\] you need to solve \[3x - 5 = 0\] for the solution...

OpenStudy (campbell_st):

can you do that..?

OpenStudy (anonymous):

x=1.67

OpenStudy (campbell_st):

correct... so the steps.. 1. factorise 2. solve the factors = 0 well done

OpenStudy (anonymous):

You factor the equation first... the factor is (3x-5) ^2. Forgot how to solve with squared factors, but you have to set it to equal 0